OpenStudy (anonymous):

Solve for x: 5x + 2 = 4x - 9 Answer 11 7 -7 -11

OpenStudy (anonymous):

-11

OpenStudy (anonymous):

5x + 2 = 4x - 9 5x - 4x = -9 - 2 x = -11